OpenStudy (anonymous):

Solve 2x - 11 - 8x = -5. -3 -1 1 3

OpenStudy (ajprincess):

Vat u get ven u add 2x and -8x?

OpenStudy (anonymous):

wattt

OpenStudy (ajprincess):

what is 2x-8x?

OpenStudy (anonymous):

-6

OpenStudy (karatechopper):

To solve: 2x - 11 - 8x = -5 We need to combine the like terms. We will combine 2x and 8x together. Making... 2x - 11 - 8x = -5 Become... -6x-11=-5 Now our new equation is.. -6x-11=-5 Now we will add 11 to both sides. -6x-11=-5 +11 +11 ------------ -6x=6 Now we can divide -6 on both sides.. So... -6x/-6=6/-6 x=-1

OpenStudy (anonymous):

thank u karatechopper

OpenStudy (karatechopper):

@ajprincess Sorry i never saw u answering. Very sorry i disturbed your teaching.:( Forgive?

OpenStudy (ajprincess):

it's k. No worries. @karatechopper

OpenStudy (karatechopper):

Thank you:)

OpenStudy (ajprincess):

yw.:D